# CaliDrop: KV Cache Compression with Calibration

### v1

@@ 5 Discussion
CaliDrop has two key hyper-parameters: $\theta_{1}$ and $\theta_{2}$. $\theta_{1}$ controls the frequency of recomputations, while $\theta_{2}$ determines the conditions under which calibration is deemed acceptable.
These hyper-parameters directly influence the model’s throughput and accuracy. Throughput is primarily affected by $\theta_{1}$, as it is directly tied to the frequency of recomputations. Accuracy, however, is influenced by both $\theta_{1}$ and $\theta_{2}$. $\theta_{1}$ affects calibration quality, with more frequent recomputations leading to better calibration. $\theta_{2}$ determines the criteria for when calibration is considered acceptable. If the conditions are too strict, the model may fail to fully leverage the benefits of calibration. On the other hand, if the conditions are too lenient, the model may suffer from low-quality calibration, which could reduce accuracy. Therefore, in this section, we focus on how to improve accuracy while minimizing the decrease in model throughput.
@@ 5.2 Speed Analysis
Figure 3: (a) Accuracy and throughput under different calibration sizes. No cali: without calibration. Full cali: all evicted tokens are used for calibration (calibration size = $\infty$). We set $\theta_{1}=0.7$ in our experiment. (b) Throughput of CaliDrop under different input lengths and $\theta_{1}$. (c) Recomputation frequency of CaliDrop in different layers under different $\theta_{1}$.
The speed of CaliDrop is primarily influenced by $\theta_{1}$. To explore this, we conduct experiments using LLaMA-3-8B-Instruct with SnapKV and a KV size of 256. We fix the batch size at 4 and set the output length to 128. In the experiments, we vary the input length and $\theta_{1}$ while monitoring throughput. The results are presented in Figure 3(a). As anticipated, a higher $\theta_{1}$ leads to more frequent recomputations, which reduces throughput. Additionally, as the input length increases, CaliDrop’s throughput slows down due to the greater demand for data transfer and computation.
In addition, we also investigate the frequency of recomputation. We conduct experiments on LongBench using LLaMA-3-8B-Instruct with SnapKV and a KV size of 128. We compare the recompute frequency (the ratio of recomputations to total decoding steps) in different layers under different $\theta_{1}$ in Figure 3(b). We find that the degree of variation in query similarity between different layers is inconsistent, leading to differences in recompute frequencies across layers. Additionally, as observed earlier, a higher $\theta_{1}$ leads to more recomputations. Overall, in our main experiment ($\theta_{1}$=0.7), recomputation occurs approximately every eight steps.
@@ 5.3 Efficiency Optimization
In the main experiment, we have not applied any efficiency optimizations. There are many ways to improve efficiency. In theory, most of the existing KV optimization methods can be used for the offloaded KV cache, such as quantization, token eviction, etc. Other methods like specialized CUDA kernel, page-level I/O (quest, ), communication overlapping is also useful for CaliDrop. We propose a promising method and demonstrate that its theoretical computation is not high (approximately every 8 steps for recalculation), and does not focus on its actual efficiency. Therefore, more efficiency optimization work will be left to future work. We here propose a simple optimization method that better achieves a balance between efficiency and accuracy. During the compression phase, when evicted tokens are offloaded, we introduce a hyper-parameter called the calibration size, which specifies the maximum number of tokens that can be offloaded. We rank the tokens by importance and offload the most important tokens (up to the calibration size) to optimize the efficiency of calibration computation. Figure 3(c) shows the accuracy of LLaMA-3-8B-Instruct using SnapKV and CaliDrop on LongBench under different calibration sizes, along with throughput results from the experiment setup in Section 5.2. We find that a small calibration size can lead to significant performance improvements with only a modest sacrifice in throughput.

### v2

@@ 5 Discussion
CaliDrop has two key hyperparameters: $\theta_{1}$ and $\theta_{2}$. Specifically, $\theta_{1}$ controls the frequency of recomputations, while $\theta_{2}$ determines the threshold conditions under which calibration is deemed acceptable.
These hyperparameters directly influence the model throughput and accuracy. Throughput is primarily affected by $\theta_{1}$, as it is directly tied to the frequency of recomputations. However, accuracy is influenced by both $\theta_{1}$ and $\theta_{2}$. $\theta_{1}$ affects calibration quality, with more frequent recomputations leading to better calibration. $\theta_{2}$ determines the criteria for when calibration is considered acceptable. If the conditions are too strict, the model may not fully utilize the benefits of calibration. On the other hand, if the conditions are too lenient, the model may suffer from low-quality calibration, which could reduce accuracy. Therefore, in this section, we focus on how to improve accuracy while minimizing the decrease in model throughput.
@@ 5.1 Influence on Accuracy
Figure 3: (a) Throughput of CaliDrop and SnapKV under different batch size. (b) Recomputation frequency of CaliDrop in different layers under different $\theta_{1}$.(c) Accuracy and throughput under different calibration sizes. No cali: without calibration. Full cali: all evicted tokens are used for calibration (calibration size = $\infty$). We set $\theta_{1}=0.7$ in our experiment.
5.2 Speed Analysis
We evaluate throughput across varying batch sizes using an input length of 1024, output length of 128, and a KV cache budget of 128 on an 80GB A800. As shown in Figure 3(a), SnapKV achieves notable throughput gains over FullKV due to its reduced memory footprint. CaliDrop adds only marginal computational overhead compared to SnapKV, while maintaining high throughput across practical batch sizes. Notably, its calibration-based accuracy improvement incurs minimal latency impact, showing that CaliDrop effectively balances efficiency and performance.
We also investigate the frequency of recomputation. We conduct experiments on LongBench using LLaMA-3-8B-Instruct with SnapKV and a KV size of 128. We compare the recompute frequency (the ratio of recomputations to total decoding steps) in different layers under different $\theta_{1}$ in Figure 3(b). We find that the degree of variation in query similarity between different layers is inconsistent, leading to differences in recompute frequencies across layers. Additionally, as observed earlier, a higher $\theta_{1}$ leads to more recomputations. Overall, in our main experiment ($\theta_{1}$=0.7), recomputation occurs approximately every eight steps.
@@ 5.3 Efficiency Optimization
In our main experiment, we did not apply any efficiency optimizations. However, various techniques can be used to improve efficiency. In principle, most existing KV cache optimization methods—such as quantization and token eviction—are applicable to the offloaded cache. Additional strategies like specialized CUDA kernels, page-level I/O (quest, ), and communication overlapping may also benefit CaliDrop. Since the theoretical computation overhead is low (recomputation occurs approximately every 8 steps), we leave extensive efficiency optimization to future work. As a lightweight yet effective enhancement, we introduce a hyper-parameter called calibration size, which controls the number of offloaded tokens by ranking them according to importance. This approach significantly reduces calibration overhead while preserving most of the accuracy benefits. Figure 3(c) shows the accuracy of LLaMA-3-8B-Instruct using SnapKV and CaliDrop on LongBench under different calibration sizes, along with throughput results from the experiment setup in Section 5.2. We find that a small calibration size can lead to significant performance improvements with only a modest sacrifice in throughput.